how would u find the surface area & volume of a square-based pyramid. the heght is 6cm and base is 4cm.
Answer by rapaljer(4671)

There are four sides that are 6cm by 4 cm = 4*6*4= 96 sq.cm.
Add to this the top and bottom which are 4 by 4, for a total of 2*16 = 32 sq.cm.

Total surface area = 96 + 32 = 128 sq.cm.

The Volume = LWH= 4*4*6= 96 cubic cm.

For additional help with Perimeters, Areas, and Volumes, please seee my own website. Click on my tutor name "rapaljer" anywhere in algebra.com. There you will see the link to my Homepage. From my Homepage, near the bottom of the page, look for the link "Topics from Geometry." There are several links that may be helpful to you. I hope this is helpful.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus